# On the Topology Awareness and Generalization Performance of Graph Neural Networks

### v1

@@ Abstract
Many computer vision and machine learning problems are modelled as learning tasks on graphs, where graph neural networks (GNNs) have emerged as a dominant tool for learning representations of graph-structured data. A key feature of GNNs is their use of graph structures as input, enabling them to exploit the graphs’ inherent topological properties—known as the topology awareness of GNNs. Despite the empirical successes of GNNs, the influence of topology awareness on generalization performance remains unexplored, particularly for node-level tasks that diverge from the assumption of data being independent and identically distributed (I.I.D.). The precise definition and characterization of the topology awareness of GNNs, especially concerning different topological features, are still unclear. This paper introduces a comprehensive framework to characterize the topology awareness of GNNs across any topological feature. Using this framework, we investigate the effects of topology awareness on GNN generalization performance. Contrary to the prevailing belief that enhancing the topology awareness of GNNs is always advantageous, our analysis reveals a critical insight: improving the topology awareness of GNNs may inadvertently lead to unfair generalization across structural groups, which might not be desired in some scenarios. Additionally, we conduct a case study using the intrinsic graph metric, the shortest-path distance, on various benchmark datasets. The empirical results of this case study confirm our theoretical insights. Moreover, we demonstrate the practical applicability of our framework by using it to tackle the cold start problem in graph active learning.
@@ 1 Introduction
Many problems in computer vision and machine learning are modeled as learning tasks on graphs. For example, in semantic segmentation, graphs model the relationships between different image regions, enhancing accuracy and context-aware segmentation. Graph neural networks (GNNs) have emerged as a dominant class of machine learning models specifically designed for learning representations of graph-structured data. They have demonstrated considerable success in addressing a wide range of graph-related problems in various domains such as chemistry [10], biology [37], social networking [6, 22], scene graph generation [51, 46] and visual relationship detection [24, 43, 49]. A defining characteristic of GNNs is their use of a spatial approach through message passing on the graph structure for feature aggregation. This enables GNNs to preserve structural information or dependencies (referred to as topology awareness) from the underlying graph structure, allowing them to be highly effective in tasks such as node classification. Fig. 1 illustrates the overall learning process of GNNs.
@@ 2.1 Topology Awareness of GNNs
Modern GNN architectures take inspiration from the Weisfeiler-Lehman isomorphism test [19, 12], which employs the graph structure to propagate information. Consequently, a significant portion of existing literature on GNNs concentrates on understanding their ability to differentiate various graph structures, known as the expressive power of GNNs [44, 28, 48, 29, 50, 30]. Inspired by the expressiveness studies, it is commonly believed that increasing topology awareness is universally beneficial and many studies focus on enabling GNNs to preserve more structural properties in the learned representation [33, 48, 29]. Nevertheless, the precise ramifications of heightened topology awareness on GNNs’ generalization performance remain shrouded in ambiguity. The intricate variations within structures across different domains and tasks further compound this complexity, underscoring the necessity for a unified framework—akin to the one proposed in this paper—to shed light on these intricacies.
@@ 2.2 Generalization of GNNs
Relatively few studies focus on understanding the generalization performance of GNNs in node-level tasks. [25, 34, 39, 3, 23, 18, 45, 41, 9] extend the conventional statistical learning or information-theoretic framework to GNNs, providing different generalization bounds based on concepts such as Vapnik–Chervonenkis dimension, mutual information, and algorithm stability. In light of this, despite empirical indications highlighting the critical role of GNNs’ topology awareness in generalization performance[48, 20, 29], there remains limited understanding of the relationship between topology awareness of GNNs and their generalization. Our proposed framework addresses this gap by utilizing the distortion concept in the approximate metric embedding to connect the topology awareness of GNNs and their generalization performance.
@@ 3.3 Topology Awareness
In the context of GNNs, evaluating distortion involves a straightforward comparison of embeddings in the embedding domain with the structural measures on the graph, as demonstrated in the case study. If the distortion rate between the graph domain and the embedding domain is close to 1, the resulting embedding would largely preserve the corresponding structural information from the graph space, thereby indicating high topology awareness. Consequently, distortion can be employed to measure the topology awareness of GNNs concerning the corresponding structural properties (e.g., shortest-path distance) measured in the graph domain. Consequently, distortion can be employed to measure the topology awareness of GNNs concerning the corresponding structural properties (e.g., shortest-path distance) measured in the graph domain. This concept of distortion is agnostic to specific structural configurations, enabling the study of diverse graph structures using appropriate distance functions. Moreover, its estimation can be readily achieved by comparing the embedding with the original structural measure, as illustrated later in the case study.

### v2

@@ Abstract
Many computer vision and machine learning problems are modelled as learning tasks on graphs, where graph neural networks (GNNs) have emerged as a dominant tool for learning representations of graph-structured data. A key feature of GNNs is their use of graph structures as input, enabling them to exploit the graphs’ inherent topological properties—known as the topology awareness of GNNs. Despite the empirical successes of GNNs, the influence of topology awareness on generalization performance remains unexplored, particularly for node-level tasks that diverge from the assumption of data being independent and identically distributed (I.I.D.). The precise definition and characterization of the topology awareness of GNNs, especially concerning different topological features, are still unclear. This paper introduces a comprehensive framework to characterize the topology awareness of GNNs across any topological feature. Using this framework, we investigate the effects of topology awareness on GNN generalization performance. Contrary to the prevailing belief that enhancing the topology awareness of GNNs is always advantageous, our analysis reveals a critical insight: improving the topology awareness of GNNs may inadvertently lead to unfair generalization across structural groups, which might not be desired in some scenarios. Additionally, we conduct a case study using the intrinsic graph metric, the shortest-path distance, on various benchmark datasets. The empirical results of this case study confirm our theoretical insights. Moreover, we demonstrate the practical applicability of our framework by using it to tackle the cold start problem in graph active learning. footnotetext: $\dagger$: corresponding authors footnotetext: To appear as a conference paper at ECCV 2024
@@ 1 Introduction
Many problems in computer vision and machine learning are modeled as learning tasks on graphs. For example, in semantic segmentation, graphs model the relationships between different image regions, enhancing accuracy and context-aware segmentation. Graph neural networks (GNNs) have emerged as a dominant class of machine learning models specifically designed for learning representations of graph-structured data. They have demonstrated considerable success in addressing a wide range of graph-related problems in various domains such as chemistry [10], biology [42], social networking [6, 36, 22], scene graph generation [55, 50] and visual relationship detection [24, 47, 53]. A defining characteristic of GNNs is their use of a spatial approach through message passing on the graph structure for feature aggregation. This enables GNNs to preserve structural information or dependencies (referred to as topology awareness) from the underlying graph structure, allowing them to be highly effective in tasks such as node classification. Fig. 1 illustrates the overall learning process of GNNs.
@@ 2.1 Topology Awareness of GNNs
Modern GNN architectures take inspiration from the Weisfeiler-Lehman isomorphism test [19, 12], which employs the graph structure to propagate information. Consequently, a significant portion of existing literature on GNNs concentrates on understanding their ability to differentiate various graph structures, known as the expressive power of GNNs [48, 28, 52, 29, 54, 30]. Inspired by the expressiveness studies, it is commonly believed that increasing topology awareness is universally beneficial and many studies focus on enabling GNNs to preserve more structural properties [38, 39] in the learned representation [33, 52, 29]. Nevertheless, the precise ramifications of heightened topology awareness on GNNs’ generalization performance remain shrouded in ambiguity. The intricate variations within structures across different domains and tasks further compound this complexity, underscoring the necessity for a unified framework—akin to the one proposed in this paper—to shed light on these intricacies.
@@ 2.2 Generalization of GNNs
Relatively few studies focus on understanding the generalization performance of GNNs in node-level tasks. [25, 34, 41, 3, 23, 18, 49, 45, 40, 9] extend the conventional statistical learning or information-theoretic framework to GNNs, providing different generalization bounds based on concepts such as Vapnik–Chervonenkis dimension, mutual information, and algorithm stability. In light of this, despite empirical indications highlighting the critical role of GNNs’ topology awareness in generalization performance[52, 20, 29], there remains limited understanding of the relationship between topology awareness of GNNs and their generalization. Our proposed framework addresses this gap by utilizing the distortion concept in the approximate metric embedding to connect the topology awareness of GNNs and their generalization performance.
@@ 3.3 Topology Awareness
In the context of GNNs, evaluating distortion involves a straightforward comparison of embeddings in the embedding domain with the structural measures on the graph, as demonstrated in the case study. If the distortion rate between the graph domain and the embedding domain is close to 1, the resulting embedding would largely preserve the corresponding structural information from the graph space, thereby indicating high topology awareness. Consequently, distortion can be employed to measure the topology awareness of GNNs concerning the corresponding structural properties (e.g., shortest-path distance) measured in the graph domain. This concept of distortion is agnostic to specific structural configurations, enabling the study of diverse graph structures using appropriate distance functions. Moreover, its estimation can be readily achieved by comparing the embedding with the original structural measure, as illustrated later in the case study.
@@ 6.1 Limitation and Future Works
Acknowledgement. We would like to thank the anonymous reviewers for their helpful comments. JS and CW are supported by grants from Hong Kong RGC under the contracts HKU 17207621, 17203522 and C7004-22G (CRF).
Appendix 0.A Proof of Theorem 4.1
In this appendix, we provide proof for Theorem 4.1.
Proof
Let $\mathcal{G}=(\mathcal{V},\mathcal{E},\mathbf{X})$ be a given graph and $\mathcal{V}_{0}\subset\mathcal{V}$ be the training set. Let $s$ be the structure of interest with distance measure $d_{s}$ and $\mathcal{V}_{i}$ be an arbitrary subgroup. Let $\mathcal{M}$ be a given GNN model with distortion $\alpha$ and scaling factor $r$ and its prediction function $g$.
Let $\sigma:\mathcal{V}_{i}\mapsto\mathcal{V}_{0}$ denote a mapping that map a vertex $v$ from $\mathcal{V}_{i}$ to the closest vertex in $\mathcal{V}_{0}$, i.e.,
$\sigma(v)=\arg\min_{u\in\mathcal{V}_{0}}d_{s}(u,v).$ (7)
Let’s consider the loss of vertex $v$, $\mathcal{L}(g\circ\mathcal{M}(v),y_{v})$. By assumption, $\mathcal{L}$ is smooth and let’s denote $B_{\mathcal{L}}^{up1}$ to be the upper bound for the first derivative with respect to $g\circ\mathcal{M}(v)$ and $B_{\mathcal{L}}^{up2}$ to be the upper bound for the first derivative with respect to $y_{v}$. Then, consider the Taylor expansion of $\mathcal{L}(g\circ\mathcal{M}(v),y_{v})$ with respect to $\sigma(v)$, which is given as follows
$\begin{split}\mathcal{L}(g\circ\mathcal{M}(v),y_{v})\leq\mathcal{L}(g\circ\mathcal{M}(\sigma(v)),y_{\sigma(v)})+B_{\mathcal{L}}^{up1}||g\circ\mathcal{M}(v)-g\circ\mathcal{M}(\sigma(v))||+B_{\mathcal{L}}^{up2}||y_{v}-y_{\sigma(v)}||\end{split}$ (8)
Next, let’s examine the inequality above term by term and start with $||g\circ\mathcal{M}(v)-g\circ\mathcal{M}(\sigma(v))||$. Let’s denote $h_{v}$ and $h_{\sigma(v)}$ the embedding for vertex $v$ and $\sigma(v)$. Then, we have that,
$||g\circ\mathcal{M}(v)-g\circ\mathcal{M}(\sigma(v))||=||g(h_{v})-g(h_{\sigma(v)})||$ (9)
By definition of distortion as given in Def. 3, we have that,
$||h_{v}-h_{\sigma(v)}||\leq r\alpha d_{s}(v,\sigma(v))$ (10)
By assumption, the prediction function $g$ is smooth. Let $B_{g}$ denote the upper bound of first derivative of the prediction function $g$. Then, we have
$||g(h_{v})-g(h_{\sigma(v)})||\leq B_{g}^{up}||h_{v}-h_{\sigma(v)}||$ (11)
Substitute all these back, we have
$||g\circ\mathcal{M}(v)-g\circ\mathcal{M}(\sigma(v))||\leq B_{g}^{up}r\alpha d_{s}(v,\sigma(v))$ (12)
Next, let’s consider $||y_{v}-y_{\sigma(v)}||$. Similarly, by the data smoothness assumption and distortion, we have that
$||y_{v}-y_{\sigma(v)}||\leq B_{l}||h_{v}-h_{\sigma(v)}||\leq B_{l}^{up}r\alpha d_{s}(v,\sigma(v))$ (13)
Substitute these back to the inequality we start with, we have that
$\begin{split}\mathcal{L}(g\circ\mathcal{M}(v),y_{v})&\leq\mathcal{L}(g\circ\mathcal{M}(\sigma(v)),y_{\sigma(v)})+B_{\mathcal{L}}^{up1}B_{g}^{up}r\alpha d_{s}(v,\sigma(v))+B_{\mathcal{L}}^{up2}B_{l}^{up}r\alpha d_{s}(v,\sigma(v))\\ &\leq\mathcal{L}(g\circ\mathcal{M}(\sigma(v)),y_{\sigma(v)})+(B_{l}^{up}B_{\mathcal{L}}^{up1}+B_{g}^{up}B_{\mathcal{L}}^{up2})r\alpha\max_{v^{\prime}\in\mathcal{V}_{i}}d_{s}(v^{\prime},\sigma(v^{\prime}))\\ &=\mathcal{L}(g\circ\mathcal{M}(\sigma(v)),y_{\sigma(v)})+(B_{l}^{up}B_{\mathcal{L}}^{up1}+B_{g}^{l}B_{\mathcal{L}}^{up2})r\alpha D_{s}(\mathcal{V}_{i},\mathcal{V}_{0}).\end{split}$ (14)
Since the inequality above holds for every vertex $v$, then we have
$\begin{split}R^{\mathcal{L}}(g\circ\mathcal{M},\mathcal{V}_{i})&\leq R^{\mathcal{L}}(g\circ\mathcal{M},\mathcal{V}_{0})+(B_{l}^{up}B_{\mathcal{L}}^{up1}+B_{g}^{up}B_{\mathcal{L}}^{up2})r\alpha D_{s}(\mathcal{V}_{i},\mathcal{V}_{0})\\ &=R^{\mathcal{L}}(g\circ\mathcal{M},\mathcal{V}_{0})+\mathcal{O}(\alpha D_{s}(\mathcal{V}_{i},\mathcal{V}_{0}))\end{split}$ (15)
Appendix 0.B Proof of Theorem 4.2
Before diving into the detailed proof, we present an outline of the structure of the proof and prove a lemma which we use in the proof of the theorem.
Outline of the proof for the thereom
  1. 1.
    Suppose we are given $\mathcal{V}_{i}$ and $\mathcal{V}_{j}$, two test groups which satisfy the premise of the theorem;
  2. 2.
    Then, we can approximate and bound the loss of each vertex in these groups based on the nearest vertex in the training set by extending the result from Theorem 4.1;
  3. 3.
    If we can show that there exists a constant independent of the property of each test group, then we obtain the results of Theorem 4.2.
Proof
Let $\mathcal{G}=(\mathcal{V},\mathcal{E},\mathbf{X})$ be a given graph and $\mathcal{V}_{0}\subset\mathcal{V}$ be the training set. Let $s$ be the structure of interest with distance measure $d_{s}$ and $\mathcal{V}_{i}$ be an arbitrary subgroup. Let $\mathcal{M}$ be a given GNN model with distortion $\alpha$ and scaling factor $r$ and its prediction function $g$.
Let $\mathcal{V}_{i}$ and $\mathcal{V}_{j}$ be the two test groups satisfy the premise of the theorem.
Let $Q_{i}:\mathcal{V}_{i}\mapsto\mathcal{V}_{0}$ be the mapping that maps vertex $u\in\mathcal{V}_{i}$ to the closest vertex $v\in\mathcal{V}_{0}$. For simplicity, we denote the vertex as $q_{u}=Q_{i}(u)$. We can do a linear approximation of $u$ around $q_{u}$ and obtain a similar upper bound in Theorem 4.1, which is given as follows:
$\begin{split}R^{\mathcal{L}}(g\circ\mathcal{M},\mathcal{V}_{i})\leq R^{\mathcal{L}}(g\circ\mathcal{M},\mathcal{V}_{0})+(B_{l}^{up}B_{\mathcal{L}}^{up1}+B_{g}^{up}B_{\mathcal{L}}^{up2})r\alpha D_{s}(\mathcal{V}_{i},\mathcal{V}_{0})\end{split}$ (16)
Similarly, let $Q_{j}:\mathcal{V}_{j}\mapsto\mathcal{V}_{0}$ be the mapping that maps vertex $u\in\mathcal{V}_{j}$ to the closest vertex $u\in\mathcal{V}_{0}$. For simplicity, we denote the vertex as $q^{\prime}_{u}=Q_{j}(u)$. Because we assume the training loss is arbitrarily small (easily obtainable with sufficient parameters) and the groups are in the local neighbourhood of the training set, we can do a linear approximation of $u$ around $q^{\prime}_{u}$ and have the following lower bound:
$\begin{split}\mathcal{L}(g\circ\mathcal{M}(v),y_{v})&\geq\mathcal{L}(g\circ\mathcal{M}(q_{v}),y_{q_{v}})+B_{\mathcal{L}}^{l1}||g\circ\mathcal{M}(v)-g\circ\mathcal{M}(q_{v})||+B_{\mathcal{L}}^{l2}{\mathcal{L}}||y_{v}-y_{q_{v}}||\\ &\geq\mathcal{L}(g\circ\mathcal{M}(q_{v}),y_{q_{v}})+(B_{l}^{l}B_{\mathcal{L}}^{l1}+B_{g}^{l}B_{\mathcal{L}}^{l2})rD_{s}(\mathcal{V}_{i},\mathcal{V}_{0}).\end{split}$ (17)
where $B_{\mathcal{L}}^{l1}$ and $B_{\mathcal{L}}^{l1}$ are the lower bound of the $\mathcal{L}(a,b)$ function with respect to $a$ and $b$. $B_{l}^{l}$ and $B_{g}^{l}$ are the lower bound for the prediction function and data generation. $B_{l1}$, $B_{l2}$,$B_{l}^{l}$,$B_{g}^{l}$ exists because $\mathcal{L},g$ and the data generation are smooth. Then, we have
$\begin{split}R^{\mathcal{L}}(g\circ\mathcal{M},\mathcal{V}_{j})&\geq R^{\mathcal{L}}(g\circ\mathcal{M},\mathcal{V}_{0})+(B_{l}^{l}B_{\mathcal{L}}^{l1}+B_{g}^{l}B_{\mathcal{L}}^{l2})rD_{s}(\mathcal{V}_{i},\mathcal{V}_{0}).\end{split}$ (18)
Now suppose $D_{s}(\mathcal{V}_{i},\mathcal{V}_{0})>D_{s}(\mathcal{V}_{j},\mathcal{V}_{0})$. Let’s consider the difference $\mathcal{V}_{i}$ and $\mathcal{V}_{j}$ with Eq. (18) being applied on $\mathcal{V}_{i}$ and Eq. (16) being applied on $\mathcal{V}_{j}$, then we have that
$\begin{split}R^{\mathcal{L}}(g\circ\mathcal{M},\mathcal{V}_{i})-R^{\mathcal{L}}(g\circ\mathcal{M},\mathcal{V}_{j})&\geq R^{\mathcal{L}}(g\circ\mathcal{M},\mathcal{V}_{0})+(B_{l}^{l}B_{\mathcal{L}}^{l1}+B_{g}^{l}B_{\mathcal{L}}^{l2})rD_{s}(\mathcal{V}_{i},\mathcal{V}_{0})-\\ &[R^{\mathcal{L}}(g\circ\mathcal{M},\mathcal{V}_{0})+(B_{l}^{up}B_{\mathcal{L}}^{up1}+B_{g}^{up}B_{\mathcal{L}}^{up2}r\alpha D_{s}(\mathcal{V}_{j},\mathcal{V}_{0})]\end{split}$ (19)
Then, we can derive that if we have
$\begin{split}(B_{l}^{l}B_{\mathcal{L}}^{l1}+B_{g}^{l}B_{\mathcal{L}}^{l2})rD_{s}(\mathcal{V}_{i},\mathcal{V}_{0})>B_{l}^{up}B_{\mathcal{L}}^{up1}+B_{g}^{up}B_{\mathcal{L}}^{up2}D_{s}r\alpha(\mathcal{V}_{j},\mathcal{V}_{0})\end{split}$ (20)
or equivalently,
$\begin{split}D_{s}(\mathcal{V}_{i},\mathcal{V}_{0})>\frac{B_{l}^{up}B_{\mathcal{L}}^{up1}+B_{g}^{up}B_{\mathcal{L}}^{up2}}{(B_{l}^{l}B_{\mathcal{L}}^{l1}+B_{g}^{l}B_{\mathcal{L}}^{l2})}\alpha D_{s}(\mathcal{V}_{j},\mathcal{V}_{0})\end{split}$ (21)
Let $\delta=\frac{B_{l}^{up}B_{\mathcal{L}}^{up1}+B_{g}^{up}B_{\mathcal{L}}^{up2}}{(B_{l}^{l}B_{\mathcal{L}}^{l1}+B_{g}^{l}B_{\mathcal{L}}^{l2})}$, we get the result desired by the theorem.
Appendix 0.C Procedure for Solving Eq. (6)
First recall that the optimization problem we want to solve is as follows.
$\min_{\mathcal{V}_{0}\subset\mathcal{V}}D_{s}(\mathcal{V}\backslash\mathcal{V}_{0},\mathcal{V}_{0})~{}~{}~{}s.t.~{}~{}~{}|\mathcal{V}_{0}|=k$ (22)
where $k$ is the given initial set size and $D_{s}(.)$ is given in Def. 2 with shortest path distance as the metric. Eq. 6 amounts to the well-known k-center problem [14] in graph theory and is NP-hard. For our experiment, we use Eq. 6 as guidance and modify the simple greedy algorithm, farthest-first traversal, to be a sampling method for obtaining a solution. We now describe how we modified the standard greedy algorithm into a sampling algorithm. We start with describing the standard greedy algorithm.
Algorithm 1 k-center Greedy
0:  $\mathcal{V}$ //vertex set of a given graph
0:  $k$ //number of centersInitialize an empty array $V$
  1. 1.
    randomly picks a random vertex $v$ to $V$
  2. 2.
    pick the next vertex to be the vertex furthest away from $V$
  3. 3.
    repeat until $|V|=k$
The procedure described above is a standard 2-approximation greedy algorithm for k-center problem. Next, we describe its sampling variant which use the distance as the chance of being sampled.
Algorithm 2 Coverage-based Sampling
0:  $\mathcal{V}$ //vertex set of a given graph
0:  $k$ //size of set to be sampledSample/select the set of size $k$ with the following procedures:
  1. 1.
    initialize $V$ with the vertex of the highest degree.
  2. 2.
    sample $u$ from $\mathcal{V}\backslash V$ based on probability $p_{u}$ proportion to $D_{s}(u,\mathcal{V}\backslash V)$, following the formula,
    $p_{u}=\frac{D_{s}(u,V)}{\sum_{v\in\mathcal{V}\backslash V}D_{s}(V,V)}$
  3. 3.
    repeat until $k$ vertices are sampled
0.C.1 Running Time Complexity
It is obvious that the procedure above runs in $\mathcal{O}(k)$ if the distance between any pair of vertices is given and runs in $\mathcal{O}(kn)$ if the distance needed to be computed for each step, where $n$ is the number of vertice.
Appendix 0.D Additional Experiments Details and Results
In this appendix, we provide additional experimental results and include a detailed set-up of the experiments for reproducibility.
0.D.1 Hardware and Software
All the experiments of this paper are conducted on the following machine
CPU: two Intel Xeon Gold 6230 2.1G, 20C/40T, 10.4GT/s, 27.5M Cache, Turbo, HT (125W) DDR4-2933
GPU: four NVIDIA Tesla V100 SXM2 32G GPU Accelerator for NV Link
Memory: 256GB (8 x 32GB) RDIMM, 3200MT/s, Dual Rank
OS: Ubuntu 18.04LTS
0.D.2 Training Hyper-parameters and Detailed Dataset Description
Table 3 gives a more detailed description of the statistic of datasets used in the experiment as well as the hyper-parameters used for training. We keep the hyper-parameters the same across different models. To keep the experiment consistent, we do not employ randomized regularization techniques such as dropout. To ensure a fair comparison, we make sure each experiment is trained till convergence.
Table 3: Description on the datasets and hyper-parameter different tasks.
Datasets Cora Citeseer Cora Full Coauthor CS OGB-Arxiv
Number of Vertexes 2,708 3,327 19,793 18,333 169,343
Number of Edges 10,556 9,228 126,842 163,788 1,166,243
Dimension of Node Feature 1,433 3,703 8,710 6,805 128
Number of Class 7 6 70 15 40
Number of Hidden Layer 2 2 2 2 2
Hidden Layer Dimension 128 128 256 256 256
Prediction Head MLP
Prediction Head Activation ReLU
Prediction Head Dimension 32 32 128 128 256
Optimizer Adam
Learning Rate 0.01
Weight Decay 0.0005
Dropout 0
Epoch 100
0.D.3 Additional Experiment Results
In addition to the datasets we present in the main text, we also conduct experiments on Citeseer, Cora Full and OGBN-arxiv. The results of these experiments show that the theoretical structures proved in this paper are indeed consistent across different datasets. Next, we present additional experimental results on these datasets for validating the main theorems of our papers.
0.D.3.1 Graph Distance and Performance
We present additional results with Citeseer, CoraFull, and OGBN-Arxiv on the relation between the graph distance (shortest-path distance). For these experiments, we follow the same setting as the one presented in Fig. 4. We compute the average prediction accuracy in each group of vertices (corresponding to vertices at each hop of the neighbourhood of the training set). ’0 hop’ indicates the training set. We observe that GNNs tend to predict better on the vertices that are closer to the training set. The results are presented in Fig. 6,  7 and  8.
0.D.3.2 Subgroup Embedding Distance and Accuracy
We present additional results with Citeseer, CoraFull, and OGBN-Arxiv on that GNNs preserve distances (with a small distortion rate) locally, by evaluating the relation between the graph distances of vertices to the training set and the embedding distances (Euclidean distance) of vertex representations to the representations of the training set, i.e., the relation between $d(v,\mathcal{D})$ and $d(\mathcal{M}_{\theta_{\mathcal{D}}}(v),\mathcal{M}_{\theta_{\mathcal{D}}}(\mathcal{D}))$. In this experiment, we follow the same setting as the one in Fig. 3 and train the GNN models on the default training set provided in the dataset. We extract vertices within the 5-hop neighbourhood of the training set and group the vertices based on their distances to the training set. We then compute the average embedding distance of each group to the training set in the embedding space. Based on Def. 3, if an embedding has low distortion, we should expect to see linear relation with slop being the scaling factor. As shown in Fig. 9,Fig. 10, and Fig. 11 there exists a strong correlation between the graph distance and the embedding distance. The relation is near linear and the relative orders of distances are mostly preserved within the first few hops. This implies that the distortion $\alpha$ is indeed small with different GNNs and that the common aggregation mechanism of GNNs can well capture local graph structure information (i.e., graph distances). It also indicates that the premise on the distortion rate, which we used to prove our theoretical results, can be commonly satisfied in practice.
0.D.4 Graph Distance and Performance
We further explore if the mean distance between different pairs of the training set and test group is related to the prediction performance of GNNs in the test group. In this experiment, we randomly sample 5 vertices from each class in the connected component as the training set $\mathcal{V}_{0}$. Then we evaluate the GNN model on the rest of the vertices $\mathcal{V}\backslash{V}_{0}$ in that connected component. The reason for a reduced training set size is to enable a larger variance (among different trials) on the mean distance between the training set and the rest of the vertices. when the training set is large, the difference (among different trials) in the mean distance between the training set and the rest of the vertices to the sampled training set is small because of the small-world phenomenon in real-life graph data [27]. The mean distance between the training set and the rest of the vertices (i.e., mean graph distance) is computed as $\frac{\sum_{u\in\mathcal{V_{0}}}d(u,\mathcal{D})}{|\mathcal{D^{\prime}}|}$. The results are shown in Fig. 5.
As Fig. 5, training GNNs on a training set with a smaller mean graph distance to the rest of the vertices, does generalize better, that better performance (test accuracy/loss) is achieved when the distance between the training set and the test group is smaller. This suggests that despite that we use the maximal distance as the definition for our framework, mean graph distance empirically also seems to be a valid indicator for generalization performance. Extending the framework to consider the average distance instead of the maximum distance would be an interesting future work.
Figure 5: Left: test accuracy vs. mean graph distance. Right: test loss vs. mean graph distance ( to the training set). GCN on Cora.
(a) GCN, Citeseer
(b) GraphSAGE, Citeseer
(c) GAT, Citeseer
(d) GCNII, Citeseer
Figure 6: Graph distance vs. accuracy. Additional Results on Citeseer
(a) GCN, CoralFull
(b) GraphSAGE, CoralFull
(c) GAT, CoralFull
(d) GCNII, CoralFull
Figure 7: Graph distance vs. accuracy. Additional Results on CoraFull
(a) GCN, OGB-Arxiv
(b) GraphSAGE, OGB-Arxiv
(c) GAT, OGB-Arxiv
(d) GCNII, OGB-Arxiv
Figure 8: Graph distance vs. accuracy. Additional Results on OGBN-Arxiv
(a) GCN, Citeseer
(b) GraphSAGE, Citeseer
(c) GAT, Citeseer
(d) GCNII, Citeseer
Figure 9: Graph distance vs. embedding distance. Additional Results on Citeseer
(a) GCN, CoraFull
(b) GraphSAGE, CoraFull
(c) GAT, CoraFull
(d) GCNII, CoraFull
Figure 10: Graph distance vs. embedding distance. Additional Results on CoraFull
(a) GCN, OGB-Arxiv
(b) GraphSAGE, OGB-Arxiv
(c) GAT, OGB-Arxiv
(d) GCNII, OGB-Arxiv
Figure 11: Graph distance vs. embedding distance. Additional Results on Ogbn-arxiv
Appendix 0.E Other Potential Applications
In addition to the initial dataset labelling problem discussed above, there are other potential applications of the proposed framework in other GNNs learning paradigms. Some examples include:
  1. 1.
    Fair k-shot learning: K-shot learning is a type of machine learning where the model is trained to classify new examples based on a small number of labelled examples in the training set. This type of learning is particularly useful in many situations where the amount of labelled data is limited. In addition to the active learning discussed earlier, k-shot learning can also be applied in choosing a representative experience replay buffer in incremental learning and transfer learning.
  2. 2.
    Fairness constraint: We can apply the derived results to ensure a fair predictive performance of GNNs for specified structural groups. As we have shown in the paper, the key factor of unfair predictive performance towards different structural groups is the distortion between structural distance and embedding distance. We can penalize the parameters with low distortion, which amounts to the following learning objective:
    $\mathcal{L}=\mathcal{L}_{\mathrm{train}}+\max_{\mathcal{V}_{i},\mathcal{V}_{j}\in\mathcal{V}_{1},...,\mathcal{V}_{m}}d(\mathrm{GNN}(\mathcal{V}_{i}),\mathrm{GNN}(\mathcal{V}_{j})),$ (23)
    where $\mathcal{V}_{1},...,\mathcal{V}_{m}$ are structural subgroups of interest.
  3. 3.
    GNNs Profolio: In this paper, we have focused on analyzing the performance of GNNs with respect to a single structure. However, it is also possible to extend our framework and consider multiple structures, represented by a set $S=\{s_{1},...,s_{m}\}$. The benefits of such an extension are twofold: first, it can provide a more detailed analysis of the GNN’s performance; second, the corresponding distortions $\alpha=\{\alpha_{1},...,\alpha_{m}\}$ can serve as a portfolio of the given GNN, enabling researchers and practitioners to easily determine which GNN is most suitable for a given task.
The proposed framework has promising potential for studying and being applied to the above-mentioned problems. Further research into investigating the actual application results would be an interesting future work.